# White-Basilisk: A Hybrid Model for Code Vulnerability Detection

### v1

@@ Appendix C Dataset Statistics Analysis
C.1 Implications for Model Design
@@
Appendix D Baseline Models
This section details the baseline models examined in our study. It is important to note that we did not train, finetune, or run any of these models ourselves. Instead, we collected and analyzed their reported metrics and configurations from their respective papers. For CodeT5 (CT5), CodeBERT (CB), UnixCoder (UC), StarCoder2 (SC2), and CodeGen2.5 (CG2.5), the information was sourced from the PrimeVul paper Ding et al. [2024]. For VulBERTa variants (VulBERTa-MLP and VulBERTa-CNN) and the baseline models (Baseline-BiLSTM and Baseline-TextCNN), the information was obtained from the original VulBERTa paper Hanif and Maffeis [2022]. For LineVul, the information was obtained from the original LineVul paper Fu and Tantithamthavorn [2022].
@@ Appendix D Baseline Models
CT5 Wang et al. [2021] Enc-Dec Multi-lingual code 220M
CB Feng et al. [2020] Encoder Bimodal (code + text) 125M
UC Guo et al. [2022] Encoder Cross-modal 125M
SC2 Li et al. [2023] Decoder The Stack v2 7B
CG2.5 Nijkamp et al. [2023] Decoder Code + natural lang. 7B
VulBERTa-MLP Hanif and Maffeis [2022] Encoder C/C++ code 125M
@@
Small Model Epochs ($<$7B) 10
@@
LineVul Fine-tuning Epochs 10
Appendix E Classification Head
@@
$x_{1}$ $=\text{Dropout}(\text{GELU}(W_{1}h+b_{1}))$ (12)
$x_{2}$ $=\text{Dropout}(\text{GELU}(W_{2}x_{1}+b_{2}))$
$x_{3}$ $=\text{LayerNorm}(x_{2})$
$y$ $=W_{3}x_{3}+b_{3}$
Appendix F Hyperparameter Details
F.1 Pretraining Hyperparameters
F.2 Fine-tuning Hyperparameters
Appendix G Handling Class Imbalance
A significant challenge in the development of AI classification models is the management of highly imbalanced datasets. In such scenarios, it is important to train the model to maintain its ability to detect minority classes effectively. In the context of source code vulnerability detection, we also encounter highly imbalanced classification data. To address this issue, we employ a dual approach affecting both data sampling and loss computation. This is implemented via a weighted function $w_{c}$, where $w_{c}$ represents the weight for class $c$, and $N_{c}$ denotes the number of samples in class $c$. The function is defined as follows:
$w_{c}=\frac{N}{2N_{c}}$ (13)
Based on the provided function, we implement two key components in our methodology. Firstly, we employ a Weighted Random Batch Sampler, a sampling mechanism that ensures each mini-batch contains a balanced representation of classes, thus mitigating the effects of dataset imbalance during training. Secondly, we implement a weighted loss function, modifying the standard cross-entropy loss by incorporating the class-specific weights $w_{c}$.
G.0.1 SIFT (Scale-Invariant Fine-Tuning)
We implement automated adversarial training using SIFT to improve the model’s resilience against adversarial examples. SIFT operates by introducing small perturbations to the input during training, encouraging the model to learn more robust features. In our implementation, we added a PerturbationLayer into the model architecture, which applies learnable perturbations to the input embeddings. The training process was designed to minimize both the task loss and the adversarial loss, the latter being computed as the difference between predictions on clean and perturbed inputs.
By adopting this technique, we have effectively imbued the model with a form of ’adversarial immunity’, rendering it more resilient against potential attacks or attempts to deceive its analysis. This enhanced robustness is particularly valuable in security-critical applications, where the reliability and consistency of the model’s performance are of paramount importance.
Appendix H Ablation Study: Combined Dataset Training
@@ H.2 Results and Analysis
H.3 Model Robustness Analysis
@@
H.4 Comparison with Individual Training
@@
Appendix I Ablation Study: Attention Mechanisms and Long-Range Vulnerability Detection
@@ I.1 Experimental Setup
Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from H.
We categorized sequences into four length bins for analysis:
  • Bin 1: 0-16,384 tokens (standard context length)
  • Bin 2: 16,384-32,768 tokens (extended context)
  • Bin 3: 32,768-65,536 tokens (long context)
  • Bin 4: 65,536-131,072 tokens (very long context)
For each bin, we monitored both memory consumption and model performance across all datasets, comparing our Infini-attention implementation against standard self-attention (eager implementation).
I.2 Memory Efficiency Analysis
Table 11 presents the peak memory consumption across different sequence lengths and attention implementations.
Table 11: Peak Memory Consumption (MB) by Sequence Length and Attention Type
Length Bin Infini-attention Standard Attention
Peak Reserved Peak Reserved
Bin 1 (0-16K) 1,338 1,442 32,409 39,696
Bin 2 (16K-32K) 1,654 1,956 OOM OOM
Bin 3 (32K-65K) 2,213 2,638 OOM OOM
Bin 4 (65K-131K) 3,322 3,848 OOM OOM
OOM: Out of Memory on NVIDIA A100 40GB GPU
The results demonstrate the significant memory advantages of our Infini-attention approach. Infini-attention shows near-linear memory scaling, increasing from 1.3GB to 3.3GB across bins. Standard attention requires 24x more memory for Bin 1 and fails entirely on longer sequences. While standard attention becomes infeasible beyond 16K tokens, Infini-attention successfully processes sequences up to 131K tokens with modest memory requirements.
I.3 Performance Analysis Across Sequence Lengths
Table 12 presents a comprehensive analysis of performance across different sequence lengths and datasets.
Table 12: Performance and Distribution Analysis Across Sequence Lengths
Sample Distribution Infini-attention Eager-attention
Dataset Bin Total Non-Vuln Vuln F1 Precision Accuracy F1 Precision Accuracy
BigVul Bin 1 18,853 17,747 1,106 0.943 0.946 0.993 0.937 0.967 0.993
Bin 2 7 5 2 0.800 0.667 0.857 - - -
Bin 3 4 3 1 1.000 1.000 1.000 - - -
VulDeePecker Bin 1 12,764 11,814 950 0.925 0.939 0.989 0.932 0.968 0.990
Bin 2 2 2 0 - - 1.000 - - -
Bin 3 2 2 0 - - 1.000 - - -
Bin 4 1 1 0 - - 1.000 - - -
PRIMEVUL Bin 1 25,411 24,715 696 0.233 0.208 0.952 0.240 0.237 0.958
Bin 2 18 15 3 0.250 0.200 0.667 - - -
Bin 3 1 1 0 - - 1.000 - - -
REVEAL Bin 1 2,269 2,062 207 0.474 0.416 0.888 0.468 0.445 0.898
Draper Bin 1 12,769 11,819 950 0.568 0.609 0.948 0.511 0.646 0.948
Several remarkable findings emerge from this analysis. White-Basilisk demonstrates remarkable effectiveness in detecting vulnerabilities in longer sequences (over 16K tokens), despite their relative rarity. In BigVul, the model achieves perfect detection (F1=1.000) for sequences in Bin 3. For Bin 2 sequences, it maintains strong performance (F1=0.800) despite the increased complexity. This effectiveness on longer sequences is particularly noteworthy given the increased difficulty of maintaining coherent attention over such distances. The infini-attention variant achieves comparable performance to full self-attention, with nearly identical metrics across major datasets (e.g., BigVul: 0.937 vs 0.943 F1). It maintains high precision while reducing computational complexity and demonstrates that linear attention is a viable alternative for vulnerability detection. In Bin 1, where most vulnerabilities occur, the model shows exceptional performance, with BigVul reaching F1=0.943, Accuracy=0.993 and VulDeePecker achieving F1=0.925, Accuracy=0.989. The model maintains effectiveness despite significant class imbalance, successfully detecting vulnerabilities even when they comprise only 2.74% of samples (PRIMEVUL) and maintaining balanced precision-recall trade-offs across length bins. Performance variations across datasets reveal interesting patterns. Stronger performance on BigVul and VulDeePecker suggests better handling of certain vulnerability types, while lower scores on PRIMEVUL indicate the challenge of detecting more subtle or complex vulnerabilities.
I.4 Comparative Analysis with Standard Attention
When comparing Infini-attention with standard attention (where possible), we observe. Infini-attention achieves comparable or better performance while using significantly less memory. Unlike standard attention, Infini-attention can process the full range of sequence lengths present in real codebases. The ability to handle longer sequences without performance degradation makes White-Basilisk suitable for analyzing entire codebases in a single pass.
I.5 Implications for Vulnerability Detection
This ablation study yields several important insights. The success of Infini-attention in maintaining high performance across sequence lengths validates our architectural choices. The model’s ability to handle sequences up to 131K tokens while maintaining accuracy demonstrates its practical utility for real-world applications. Strong performance on longer sequences suggests effective capture of long-range dependencies, crucial for detecting vulnerabilities that span multiple functions or files. The memory efficiency of our approach makes it feasible to deploy White-Basilisk on standard hardware, even for processing very long sequences.
These findings confirm that White-Basilisk successfully addresses the key challenges in vulnerability detection: maintaining high accuracy across varying sequence lengths while remaining computationally efficient. The model’s particular strength in handling long sequences, combined with its consistent performance on more common shorter sequences, makes it a practical and effective tool for real-world code security applications.
Appendix J Ablation Study: CWE-Specific Performance Analysis
J.1 Experimental Setup
  • Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from H.
@@ I.5 Implications for Vulnerability Detection
J.1.1 Dataset-Specific Performance Patterns

### v4

@@ Appendix D Dataset Statistics Analysis
D.1 Implications for Model Design
@@
Appendix E Baseline Models
This section details the baseline models examined in our study. It is important to note that we did not train, finetune, or run any of these models ourselves. Instead, we collected and analyzed their reported metrics and configurations from their respective papers. For CodeT5 (CT5), CodeBERT (CB), UnixCoder (UC), StarCoder2 (SC2), and CodeGen2.5 (CG2.5), the information was sourced from the PrimeVul paper Ding et al. (2024). For VulBERTa variants (VulBERTa-MLP and VulBERTa-CNN) and the baseline models (Baseline-BiLSTM and Baseline-TextCNN), the information was obtained from the original VulBERTa paper Hanif & Maffeis (2022).
@@ Appendix E Baseline Models
CT5 Wang et al. (2021) Enc-Dec Multi-lingual code 220M
CB Feng et al. (2020) Encoder Bimodal (code + text) 125M
UC Guo et al. (2022) Encoder Cross-modal 125M
SC2 Li et al. (2023) Decoder The Stack v2 7B
CG2.5 Nijkamp et al. (2023) Decoder Code + natural lang. 7B
VulBERTa-MLP Hanif & Maffeis (2022) Encoder C/C++ code 125M
@@
Small Model Epochs (<<7B) 10
@@
Appendix F Computational and Memory Complexity Analysis
This section provides a comprehensive analysis of the computational and memory complexity characteristics of our Basilisk Self-Attention mechanism, comparing it systematically against standard self-attention and the original Infini-attention approach (Munkhdalai et al., 2024).
F.1 Standard Self-Attention Process and Complexity
Standard self-attention computes attention scores for all token pairs simultaneously, requiring the following operations:
@@ F.1 Standard Self-Attention Process and Complexity
    Attention Weight Normalization: $\text{softmax}(A)$ with complexity $O(n^{2})$
@@
    Output Computation: $O=\text{softmax}(A)V$ with complexity $O(n^{2}d)$
The dominant terms are the attention matrix computation and output computation, yielding overall computational complexity of $O(n^{2}d+nd^{2})$. For typical transformer dimensions where $n>d$ for long sequences, this becomes $O(n^{2}d)$. Memory complexity is $O(n^{2})$ for storing the attention matrix, making this approach computationally prohibitive for long sequences.
F.2 Original Infini-attention Process and Complexity
The original Infini-attention processes sequences in fixed-size segments while maintaining a compressive memory state:
  1. 1.
    Segment Processing: Divide input sequence into segments of size $S$, complexity $O(1)$
  2. 2.
    Local Attention: For each segment, compute standard attention with complexity $O(S^{2}d)$
  3. 3.
    Memory Retrieval: Retrieve from compressive memory:
    where $Q$ is $(S\times d)$, $M$ is $(d\times d)$. Matrix operations yield complexity $O(Sd^{2})$
  4. 4.
    Memory Update: Update compressive state:
    $M$ $\leftarrow M+(\text{ELU}(K)+1)^{T}V$ (19)
    $z$ $\leftarrow z+\sum_{i=1}^{S}(\text{ELU}(K_{i})+1)$ (20)
    where $K^{T}$ is $(d\times S)$ and $V$ is $(S\times d)$, yielding $(d\times d)$ update with complexity $O(Sd^{2})$
  5. 5.
    Streaming Output: Process and discard segment results immediately, complexity $O(1)$
For $n/S$ segments, total computational complexity becomes: $O\left(\frac{n}{S}\times(S^{2}d+2Sd^{2})\right)=O(nSd+2nd^{2})$
Memory complexity remains constant at $O(d^{2})$ for the compressive memory state.
F.3 Basilisk Self-Attention Process and Complexity
Our Basilisk Self-Attention modifies the original Infini-attention by accumulating segment outputs for global context integration:
  1. 1.
    Segment Processing: Identical segmentation as Infini-attention with complexity $O(1)$
  2. 2.
    Dual Attention Computation: For each segment, compute both:
    • Local dot-product attention: $A_{\text{dot}}=\text{softmax}(QK^{T}/\sqrt{d})V$ with complexity $O(S^{2}d)$
    • Memory-based attention using Equation (4) with complexity $O(Sd^{2})$
  3. 3.
    Memory State Update: Update compressive memory (identical to Infini-attention) with complexity $O(Sd^{2})$:
    $M$ $\leftarrow M+(\text{ELU}(K)+1)^{T}V$ (21)
    $z$ $\leftarrow z+\sum_{i=1}^{S}(\text{ELU}(K_{i})+1)$ (22)
  4. 4.
    Output Accumulation: Store segment outputs in memory with complexity $O(Sd)$ per segment:
    total_mem_outputs $\leftarrow\text{total\_mem\_outputs}+[A_{\text{mem}}]$ (23)
    total_attn_outputs $\leftarrow\text{total\_attn\_outputs}+[A_{\text{dot}}]$ (24)
  5. 5.
    Global Integration: After processing all segments, concatenate with complexity $O(nd)$:
    total_mem $=\text{concat}(\text{total\_mem\_outputs})$ (25)
    total_attn $=\text{concat}(\text{total\_attn\_outputs})$ (26)
  6. 6.
    Adaptive Gating: Combine outputs using learnable parameter with complexity $O(nd)$:
    $O=\text{sigmoid}(\beta)\odot\text{total\_mem}+(1-\text{sigmoid}(\beta))\odot\text{total\_attn}$ (27)
For $n/S$ segments, the total computational complexity becomes: $O\left(\frac{n}{S}\times(S^{2}d+2Sd^{2}+Sd)\right)+O(nd)+O(nd)=O(nSd+2nd^{2}+2nd)$
Memory complexity is $O(nd)+O(d^{2})$ due to accumulation of all segment outputs plus the compressive memory state.
F.4 Complexity Comparison
Table 10 summarizes the complexity characteristics of all three approaches.
Table 10: Detailed Complexity Comparison of Attention Mechanisms
Mechanism Computational Memory Sequence Limit
Standard Attention $O(n^{2}d+nd^{2})$ $O(n^{2})$ Available memory
Infini-attention $O(nSd+2nd^{2})$ $O(d^{2})$ Theoretically infinite
Basilisk Attention $O(nSd+2nd^{2}+2nd)$ $O(nd)+O(d^{2})$ Available memory
Complexity Analysis: For standard attention, when processing long sequences where $n\gg d$, the quadratic term $O(n^{2}d)$ dominates. For segment-based approaches, the complexity depends on the relationship between segment size $S$, sequence length $n$, and hidden dimension $d$:
  • When $d^{2}\gg Sd$ (i.e., $d\gg S$), both segment-based methods approach $O(nd^{2})$ complexity
  • With our model parameters ($S=2048$, $d=512$), we have $S>d$, so the $nSd$ term remains significant
  • For very long sequences where $n$ is much larger than both $S$ and $d$, the $nd^{2}$ terms eventually dominate
F.5 Memory Efficiency Validation
To validate our theoretical analysis, we conducted empirical memory consumption measurements across different sequence lengths using NVIDIA A100 40GB GPU. Table 11 presents peak memory consumption for our Basilisk Self-Attention compared to standard attention.
Table 11: Empirical Memory Consumption Analysis
Length Bin Range Basilisk Attention Standard Attention
Peak (MB) Reserved (MB) Peak (MB) Reserved (MB)
Bin 1 [0, 16,384] 1,338 1,442 32,409 39,696
Bin 2 [16,384, 32,768] 1,654 1,956 OOM OOM
Bin 3 [32,768, 65,536] 2,213 2,638 OOM OOM
Bin 4 [65,536, 131,072] 3,322 3,848 OOM OOM
OOM: Out of Memory on NVIDIA A100 40GB GPU
The empirical results confirm our theoretical analysis. Basilisk Self-Attention demonstrates near-linear memory scaling, increasing from 1.3GB to 3.3GB across sequence length bins, while standard attention requires 24× more memory for sequences up to 16K tokens and becomes entirely infeasible beyond this threshold.
F.6 Practical Implications
The complexity analysis reveals several key insights:
  1. 1.
    Computational Efficiency: Our approach maintains identical computational complexity to Infini-attention while enabling more flexible global context modeling compared to the original streaming approach.
  2. 2.
    Memory Scaling: The linear memory growth $O(n\times d)$ represents a practical trade-off, allowing processing of extremely long sequences while remaining substantially more memory-efficient than quadratic attention approaches.
  3. 3.
    Implementation Flexibility: By accumulating segment outputs, our approach enables various post-processing operations and global context integration strategies that would be difficult to implement in the original streaming framework.
  4. 4.
    Sequence Length Limitations: While our approach cannot process theoretically infinite sequences like the original Infini-attention, the practical sequence length limitations are determined by available system memory rather than algorithmic constraints, making it suitable for most real-world applications.
The empirical validation demonstrates that this architectural choice enables processing of sequences that would be computationally infeasible with standard transformer architectures, while maintaining reasonable memory requirements for deployment in practical applications.
Appendix G Ablation Study: Combined Dataset Training
@@ G.2 Results and Analysis
G.3 Model Robustness Analysis
@@
G.4 Comparison with Individual Training
@@
Appendix H Ablation Study: Attention Mechanisms and Long-Range Vulnerability Detection
@@
H.1 Experimental Setup
Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from G.
We categorized sequences into 6 length bins for analysis.
H.2 Performance Analysis Across Sequence Lengths
Table 13 presents a comprehensive analysis of performance across different sequence lengths and datasets, organized into six bins based on token length ranges.
Table 13: Performance Analysis Across Sequence Length Bins
Performance Metrics Class Distribution
Dataset Bin Length Range F1 Precision Recall Accuracy Total Vuln Non-Vuln Avg Length
BigVul Bin 1 [0, 512] 0.921 0.933 0.909 0.993 15,879 692 15,187 154
Bin 2 [512, 1024] 0.983 0.975 0.990 0.996 1,829 200 1,629 709
Bin 3 [1024, 2048] 0.996 1.000 0.991 0.999 801 116 685 1,401
Bin 4 [2048, 4096] 0.957 0.930 0.985 0.976 252 67 185 2,816
Bin 5 [4096, 16384] 0.939 0.886 1.000 0.957 92 31 61 6,938
Bin 6 [16384, 131072] 0.857 0.750 1.000 0.909 11 3 8 30,581
PRIMEVUL Bin 1 [0, 512] 0.111 0.105 0.117 0.975 19,700 265 19,435 178
Bin 2 [512, 1024] 0.192 0.160 0.239 0.911 3,225 142 3,083 720
Bin 3 [1024, 2048] 0.357 0.324 0.399 0.869 1,569 143 1,426 1,416
Bin 4 [2048, 4096] 0.320 0.284 0.367 0.785 650 90 560 2,763
Bin 5 [4096, 16384] 0.411 0.341 0.518 0.689 267 56 211 7,336
Bin 6 [16384, 131072] 0.250 0.200 0.333 0.684 19 3 16 24,020
VulDeePecker Bin 1 [0, 512] 0.955 0.961 0.948 0.994 14,093 865 13,228 149
Bin 2 [512, 1024] 0.645 0.567 0.747 0.950 1,295 79 1,216 705
Bin 3 [1024, 2048] 0.387 0.545 0.300 0.956 427 20 407 1,393
Bin 4 [2048, 4096] 0.167 0.200 0.143 0.932 147 7 140 2,734
Bin 5 [4096, 16384] 0.857 0.750 1.000 0.979 48 3 45 6,040
Bin 6 [16384, 131072] 0.000 0.000 0.000 1.000 5 0 5 45,345
Draper Bin 1 [0, 512] 0.555 0.602 0.515 0.957 103,478 5,421 98,057 218
Bin 2 [512, 1024] 0.591 0.618 0.566 0.908 21,046 2,478 18,568 701
Bin 3 [1024, 2048] 0.601 0.649 0.560 0.911 2,913 350 2,563 1,215
Bin 4 [2048, 4096] 0.750 1.000 0.600 0.947 38 5 33 2,398
Bin 5 [4096, 16384] 0.000 0.000 0.000 1.000 1 0 1 4,738
Several significant findings emerge from this comprehensive analysis that provide insights into the effectiveness of our novel Basilisk self-attention mechanism across varying sequence lengths. The results reveal a compelling pattern where performance often improves with increasing sequence length, contrary to the typical expectation of degradation in long-sequence processing tasks.
The counterintuitive improvement in performance with longer sequences demonstrates that our attention mechanism effectively leverages increased contextual information. Short sequences in Bin 1, while computationally efficient to process, may lack sufficient context for the model to make optimal predictions. This is particularly evident in PRIMEVUL, where F1 scores improve from 0.111 in Bin 1 to 0.411 in Bin 5. The lower F1 score in shorter ranges likely reflects the fundamental challenge that limited context provides insufficient information for definitive classification, leading to more conservative predictions.
Most significantly, the substantial performance improvements occur precisely at the 2048-token threshold where our Basilisk self-attention mechanism becomes active. This pattern is exemplified in BigVul, where the model maintains exceptional performance through longer bins. Similarly, VulDeePecker shows remarkable recovery at Bin 5 (F1=0.857) after declining in middle ranges. This correlation between the activation threshold of our novel attention mechanism and improved performance provides empirical validation of our architectural design, demonstrating that the segmented attention approach successfully captures long-range dependencies that are crucial for complex classification tasks.
The segmented attention approach proves particularly effective for sequences exceeding 2048 tokens, where traditional attention mechanisms typically struggle with computational complexity and memory constraints. The consistent high performance across longer bins demonstrates that our mechanism successfully maintains both efficiency and effectiveness as sequence length increases. This capability is particularly valuable for applications requiring analysis of extended sequential data, whether in natural language processing, code analysis, or other domains involving long-context dependencies.
However, it is important to acknowledge the limitation imposed by the natural distribution of sequence lengths in real-world datasets. The number of samples in the highest token length bins is considerably limited—with only 11 samples in BigVul Bin 6, 19 in PRIMEVUL Bin 6, and 5 in VulDeePecker Bin 6. This sample scarcity constrains the statistical power of our validation for ultra-long sequences and limits our ability to draw definitive conclusions about model performance at the extreme end of the length spectrum. While the observed results are encouraging and demonstrate the mechanism’s potential, more extensive evaluation on larger collections of long-sequence samples would strengthen the validation of our architectural capabilities for processing sequences exceeding 16K tokens.
H.2.1 Full-File PRIMEVUL Performance Analysis
This experiment analyzed complete C/C++ source files from PRIMEVUL rather than isolated vulnerable functions. We cross-matched vulnerable samples from the validation set with their corresponding complete source files and performed classification on entire files up to 524K tokens. Note that PRIMEVUL only provides complete C/C++ files for vulnerable samples, not for clean samples, which explains the distribution patterns observed.
Table 14: Full-File PRIMEVUL Performance Across Sequence Length Bins
Bin Length Range Samples F1 Precision Recall Accuracy Vuln% Avg Length
Bin 1 [0, 16384] 23,713 0.244 0.177 0.390 0.962 1.6 524
Bin 2 [16384, 32768] 109 0.492 0.889 0.340 0.394 86.2 23,738
Bin 3 [32768, 65536] 75 0.520 1.000 0.351 0.360 98.7 45,178
Bin 4 [65536, 131072] 38 0.417 1.000 0.263 0.263 100.0 86,464
Bin 5 [131072, 262144] 11 0.167 1.000 0.091 0.091 100.0 158,538
Bin 6 [262144, 524288] 2 0.667 1.000 0.500 0.500 100.0 288,283
The data distribution reflects PRIMEVUL’s structure: Bin 1 contains both vulnerable and clean samples (1.6
F1 scores peak in Bins 2-3 (0.492-0.520) where sufficient context combines with adequate sample sizes. Perfect precision (1.000) emerges for sequences exceeding 32K tokens, but recall deteriorates with length, indicating conservative predictions on longer sequences. The small sample sizes in longer bins (11 samples in Bin 5, 2 in Bin 6) limit statistical confidence for ultra-long sequences, though maintained precision demonstrates the model’s capability to process extended context without prediction quality degradation.
Appendix I Ablation Study: CWE-Specific Performance Analysis
@@ I.1 Experimental Setup
  • Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from G.
@@
I.1.1 Dataset-Specific Performance Patterns